# Provable Low-Frequency Bias of In-Context Learning of Representations

### v1

@@ 1 Introduction
It has become a major challenge for today’s machine learning community to understand how large language models (LLMs) perform in-context learning (ICL), i.e. the ability to learn patterns or tasks solely from input sequences without any gradient updates (Brown et al., 2020; Min et al., 2022; Garg et al., 2022; Akyürek et al., 2022a). Empirical studies have demonstrated that LLMs can carry out a variety of tasks, including logical reasoning (Wei et al., 2022), programming (Gao et al., 2023), and solving mathematical problems (Hendrycks et al., 2021); recent work also suggests that LLMs are able to stay robust against noisy prompts (Cheng et al., 2025; Alazraki et al., 2025). However, the mechanisms underlying these capabilities remain largely elusive. A particularly striking phenomenon, recently highlighted by Park et al. (2024), shows that when a pre-trained LLM is fed a sequence generated by a random walk on a planar graph, with each node corresponds to a word (see “Data Generating Process” part of Figure 1), the model’s hidden representations converge to a state that reflects the original graph structure, even though the graph itself was never explicitly provided. We refer to this emergent behavior as In-Context Learning of Representations (ICLR).
@@
    Context-wise Process:  If the attention map “reflects” a function that is depends only on token identities (formally defined in Definition 2 and Definition 3), and the representations converge to a set of latent representations as the context length increases (formally defined in Definition 1), then the attention effectively applies the reflected function to the latent representations. This result is formally shown in Theorem 1;
@@
    Layer-wise Process:  The latent representations across layers, and eventually converge to a state that captures the distributional properties of the input sequence. Under the specific DGP considered in Park et al. (2024), we prove that this convergence exactly yields the representations characteristic of the ICLR phenomenon. This result is formally stated in Theorem 2.
@@
The concept of Double Convergence is illustrated in Figure 1. While our main analysis focuses on the specific DGP used in Park et al. (2024), the techniques and theoretical insights we develop are able to be extended beyond this particular setting. In Appendix D, we provide a generalized framework that is decoupled from any specific DGP, highlighting the broader applicability of our results.
@@
    We identify the double convergence process, which serves as a general framework for studying the evolution of the representation in ICL (Sections 4 and 5);
@@
    We provide theoretical explanations for several previously unexplained phenomena in Park et al. (2024), including why ICL can suppress the original meaning of each word learned in pre-training (Section 6.1), why the learned representations form an apparently regular yet slightly distorted structure (Sections 6.2 and 6.3), and why the energy decreases but does not converge to zero (Section 6.4);
@@
    Our theory highlights an implicit low-frequency bias in ICL, and predicts that LLMs are naturally robust to errors in the input prompts, which we have verified empirically (Section 6.5).
This paper is organized as follows. In Section 2, we review prior work on the mechanisms of in-context learning and the structure of hidden representations, emphasizing the core challenges and introducing core ideas of our approach. In Section 3, we define notations used and the problems considered throughout the paper, as well as outline our method. In Section 4, we present our results on context-wise convergence, and in Section 5, we establish the main result by combining context-wise and layer-wise convergence. In Section 6, we discuss the theoretical implications of our framework and address specific empirical observations. Finally, we conclude in Section 7 with a summary and outlook for future work. Due to space constraints, we defer the formal proofs of our theoretical results and the empirical validation of our assumptions to the appendix.
@@ 3 Preliminaries
Throughout this paper, we use bold upper-case letters to represent matrices (e.g. ${\bm{A}}$), bold lower-case letters to represent vectors (e.g. ${\bm{x}}$) and calligraphic upper-case letters to represent sequences of vectors (e.g. ${\mathcal{V}}=\{{\bm{v}}_{k}\}_{k=1}^{n}$). For a matrix or a vector, we use plain lower-case letters to represent their entries (e.g. $a_{i,j}$ represents the $i,j$-th entry of ${\bm{A}}$). For a number $n\in\mathbb{N}$, we denote $\{1,2,\cdots,n\}$ by $[n]$. For a set $S$, we use $\mathbbm{2}^{S}$ to represent its power set. We use ${\bm{1}}$ to represent a vector with all-entries being $1$, whose dimensionality is inferred from context. For a logical statement $\phi$, we define $\mathbbm{1}_{\{\phi\}}=\begin{cases}1&\phi\text{ is true}\\ 0&\phi\text{ is false}\end{cases}$ to be its indicator function. Given a sequence of vectors ${\mathcal{Z}}=\left\{{\bm{z}}_{x}\right\}_{x=1}^{c}\in\left(\mathbb{R}^{d}\right)^{c}$, we define ${\bm{Z}}=\mathop{\mathrm{mat}}\mathcal{Z}\in\mathbb{R}^{d\times c}$ to be the matrix formed by column-wise stacking of the vectors in $\mathcal{Z}$, i.e. the $i$-th column of ${\bm{Z}}$ is ${\bm{z}}_{i}$. Given a matrix ${\bm{A}}\in\mathbb{R}^{n\times n}$ and a sequence of vectors $\mathcal{V}=\left\{{\bm{v}}_{k}\right\}_{k=1}^{n}\in\left(\mathbb{R}^{d}\right)^{n}$, let $\bm{A}\mathcal{V}$ to be a sequence defined as ${\bm{A}}{\mathcal{V}}=\left\{\sum_{j=1}^{n}a_{k,j}{\bm{v}}_{j}\right\}_{k=1}^{n}$. For a vector function $\sigma:\mathbb{R}^{d}\to\mathbb{R}^{d}$ and a matrix ${\bm{Z}}\in\mathbb{R}^{c\times d}$, we use $\sigma({\bm{Z}})$ to represent applying $\sigma$ column-wisely to ${\bm{Z}}$.
@@ 3.1 Data Generation
Throughout this paper, we assume the input is a sequence of tokens $\mathcal{X}=\{x_{k}\}_{k=1}^{n}\in[c]^{n}$, where $n$ is the sequence length and $c$ is the vocabulary size (each token is simply a number in $[c]$). We assume that the sequence is sufficiently long, i.e., $n>10c$.
@@ 3.2 Model Architecture
Throughout the paper, we consider a simplified yet deep and non-linear Transformer model, as described in Algorithm 1, where $d\in\mathbb{N}$ is the hidden and input dimension, $L\in\mathbb{N}$ is the number of layers, ${\bm{A}}^{(\ell)}=\left\{a_{k,j}^{(\ell)}\right\}_{k,j\in[n]}$ is the (single-head) attention map at layer $\ell$, and $\sigma^{(\ell)}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the neuron-wise transformations at layer $\ell$ (which may include feedforward networks (FFNs), normalization layers, and other non-linearities).
As discussed in Section 2, the most critical simplification in our model is that we treat the attention maps as given, instead of generated from hidden representations. In exchange for this simplification, we are able to explicitly characterize the structure of the hidden representations in a deep and non-linear model, enabling us to rigorously explain multiple in-context learning behaviors. This contrasts with prior works that remain entangled in the complexity of layer-wise interactions in full Transformer architectures. Moreover, we validate our assumption empirically on real models in Appendix F, finding that our assumptions on self-attention effectively explains more than $70\%$ of actual attention connections, indicating its practical justifiability.
This model described in Algorithm 1 also omits several other standard components such as normalizations and residual connections. We note that this is to not over-complicating the theoretical results while still capturing the core mechanisms and challenges of the model. There is flexibility in our analysis to include other components, but we choose to focus on a minimal version in the main paper to clearly highlight the key ideas behind our theoretical results. See Appendix E for further discussion on integrating additional components.
@@ 3.3 Methodology Outline
As noted before, we treat the attention maps in the Transformer as given, rather than dynamically generated from the hidden representations. Specifically, we assume the attention maps in the model are composed of a class of structured maps we refer to as nice attentions, formally defined in Definition 3. These are attention maps whose connectivity patterns are determined by a function of the input tokens. This concept can be viewed as a generalization of the notion of induction heads (see eq. 1), as illustrated by the structural similarity between eq. 1 and eq. 4.
@@ 4 Context-wise Convergence
Notice that in Definition 3, the $k$-th row of the attention map $\{a_{k,j}\}_{j=1}^{n}$ is only well defined only there exists $j\in[k]$ such that $x_{j}\in f(x_{k})$ (otherwise all attention weights in this row are 00, violating the assumption that each row of ${\bm{A}}$ sums up to $1$). This is guaranteed under our setup because we have explicitly set the first $c$ tokens to be a traversal over the entire vocabulary.
We are now ready to state the main theorem of this section. Notice that this theorem stated here depends on the DGP, as it relies on the distribution of $F_{x,k}$. However, it is possible to prove a weaker but more general version of this theorem that is independent of the DGP. See Appendix D for more details.
@@
There exists a constant $C>0$ that only depends on $\mathcal{G}$, and an event with probability at least $0.999$, such that within this event, the following statement holds. Suppose ${\mathcal{V}}=\left\{{\bm{v}}_{k}\right\}_{k=1}^{n}\in\left(\mathbb{R}^{d}\right)^{n}$ is a good sequence converging to ${\mathcal{Z}}=\left\{{\bm{z}}_{x}\right\}_{x\in[c]}$ with parameter $\gamma$, and ${\bm{A}}=\left\{a_{k,j}\right\}_{k,j\in[n]}\in\mathbb{R}^{n\times n}$ is a nice attention map with parameter $\psi$ that reflects a function $f:[c]\to\mathbbm{2}^{[c]}$. Then ${\bm{A}}{\mathcal{V}}$ is a good sequence converging to
@@
Theorem 1 shows that applying a nice attention map that reflects a function to a good sequence yields another good sequence, and that the attention operation effectively acts on the latent representations. In other words, attention maps of this kind preserve convergence and transform latent representations in a token-consistent way.
@@ 5 Layer-wise Convergence
In Theorem 1, we showed how latent representations evolve under a single attention layer. In this section, we study how these latent representations change across layers.
@@
where $\rho_{\tau}^{(\ell)}\geq 0$ ($\tau\in\{{A},{B},{O},{T}\}$) is the weight of the $\tau$-th type connections, and $\sum_{\tau\in\{{A},{B},{O},{T}\}}\rho^{(\ell)}_{\tau}=1$. Empirically, we find that these four types explain over $70\%$ of attention connections in real models (see Appendix F for details), highlighting the empirical soundness of this classification.
@@ 5.1 Main Results
There exists an event with probability at least $0.99$ such that the following statement holds. Let $n_{[x]}$ be the largest $k\in[n]$ such that $x_{k}=x$. Let ${\bm{M}}={\bm{D}}^{-1/2}{\bm{W}}{\bm{D}}^{-1/2}$. Let the eigen-decomposition of ${\bm{M}}$ be
@@
where the eigenvalues are arranged in a non-increasing order of absolute values; ${\bm{\Lambda}}$ contains $q-1$ eigenvalues and ${\bm{\Lambda}}^{\prime}$ contains $n-q$ eigenvalues. Let $\bm{A}^{(\ell)}$ be defined as in eq. 7, and $\left\{{\bm{v}}_{k}^{(\ell)}\right\}_{k=1}^{n}$ be defined as in Algorithm 1. Suppose each $\sigma^{(\ell)}$ is a great mapping w.r.t. $\left(\gamma_{1}^{(\ell)},\gamma_{2}^{(\ell)},\mathbb{R}^{d},\bm{X}\right)$ and $\left(\gamma_{1}^{\prime(\ell)},\gamma_{2}^{\prime(\ell)},\mathbb{R}^{d},\bm{Y}\right)$, and have finite Lipschitz constant. Suppose there exists $\epsilon>0$ satisfying that $\delta_{q}\left(\rho_{A}^{(\ell)}{\bm{I}}+\rho_{B}^{(\ell)}{{\bm{M}}}\right)\frac{\gamma_{1}^{(\ell)}}{\gamma_{2}^{\prime(\ell)}}\leq 1-\epsilon$ for all $\ell\in[L]$. Then
@@ 6 Discussion About the Theoretical Results
In Theorem 2, we proved that the representations converge to the top eigenspace of ${\bm{M}}={\bm{D}}^{-1/2}{\bm{W}}{\bm{D}}^{-1/2}$. This matrix has been extensively studied in the literature on graph learning and spectral methods (Kipf & Welling, 2016; Li et al., 2018; Wu et al., 2019; Yang et al., 2021; Spielman, 2019; Trevisan, 2013). Specifically, since ${\bm{W}}$ is a symmetric matrix with non-negative entries, it defines a (weighted) undirected graph. Let $\widehat{\bm{L}}$ be the symmetrically normalized Laplacian matrix of this graph (Spielman, 2019), it holds that ${\bm{M}}={\bm{I}}-\widehat{\bm{L}}$, i.e. ${\bm{M}}$ and $\widehat{\bm{L}}$ share the same set of eigenvectors, with the order of eigenvalues being reversed. Thus, top eigenvectors of ${\bm{M}}$ corresponds to low eigenvalues of $\widehat{\bm{L}}$, which is known to encode the low-frequency (smooth) and low-energy signals on the graph, as they tends to assign similar values to adjacent nodes. These low eigenvectors of $\widehat{\bm{L}}$ are often used as coordinates for graph visualization, as it is know that they form figures that match human intuition (Tutte, 1963) [^2] Likely because humans also have a low-frequency bias in visual processing., and exactly explains why the ICLR phenomenon, where the hidden representations encode global graph structure, emerges in such settings.
We confirm this theoretical prediction by reproducing the experiments in Park et al. (2024) and compare the principal components of the actual hidden representations and the analytical prediction, i.e. top eigenvectors of ${\bm{W}}$. The result is shown in in Figure 2. It is clear that the analytical predictions align closely with the empirical principle components.
@@ 6.2 Why Start From the 2nd Eigenvector?
In both Theorem 2 and Figure 2, we intentionally omit the first eigenvector of ${\bm{M}}$. This is due to the coincidental alignment between the Laplacian and PCA. In short, the 1st eigenvector of ${\bm{M}}$ corresponds to a constant vector added to each ${\bm{z}}_{x}^{\prime}$; however, PCA involves a centralization step that removes the mean component from the data. Specifically, it can proved that the first eigenvector of ${\bm{M}}$ is exactly ${\bm{d}}^{-1/2}$ (Spielman, 2019), and the centralization operation is to projecting ${\bm{V}}_{n}^{(L)}$ onto the space orthogonal to ${\bm{1}}$, which is equivalent to projecting ${\bm{V}}_{n}^{(L)}{\bm{D}}^{1/2}$ onto the space orthogonal to ${\bm{d}}^{-1/2}$, effectively removing the component aligned with the first eigenvector of ${\bm{M}}$.

### v2

@@ 1 Introduction
It has become a major challenge for today’s machine learning community to understand how large language models (LLMs) perform in-context learning (ICL), i.e. the ability to learn patterns or tasks solely from input sequences without any gradient updates (Brown et al., 2020; Min et al., 2022; Garg et al., 2022; Akyürek et al., 2022a). Empirical studies have demonstrated that LLMs can carry out a variety of tasks, including logical reasoning (Wei et al., 2022), programming (Gao et al., 2023), and solving mathematical problems (Hendrycks et al., 2021); recent work also suggests that LLMs are able to stay robust against noisy prompts (Cheng et al., 2025; Alazraki et al., 2025). However, the mechanisms underlying these capabilities remain largely elusive. A particularly striking phenomenon, recently highlighted by Park et al. (2024), shows that when a pre-trained LLM is fed a sequence generated by a random walk on a planar graph, with each node corresponds to a word (see “Data Generating Process” part of Figure˜1), the model’s hidden representations converge to a state that reflects the original graph structure, even though the graph itself was never explicitly provided. We refer to this emergent behavior as In-Context Learning of Representations (ICLR).
@@
    Context-wise Process:  If the attention map “reflects” a function that is depends only on token identities (formally defined in Definition˜2 and Definition˜3), and the representations converge to a set of latent representations as the context length increases (formally defined in Definition˜1), then the attention effectively applies the reflected function to the latent representations. This result is formally shown in Theorem˜1;
@@
    Layer-wise Process:  The latent representations across layers, and eventually converge to a state that captures the distributional properties of the input sequence. Under the specific DGP considered in Park et al. (2024), we prove that this convergence exactly yields the representations characteristic of the ICLR phenomenon. This result is formally stated in Theorem˜2.
@@
The concept of Double Convergence is illustrated in Figure˜1. While our main analysis focuses on the specific DGP used in Park et al. (2024), the techniques and theoretical insights we develop are able to be extended beyond this particular setting. In Appendix˜D, we provide a generalized framework that is decoupled from any specific DGP, highlighting the broader applicability of our results.
@@
    We identify the double convergence process, which serves as a general framework for studying the evolution of the representation in ICL (Sections˜4 and 5);
@@
    We provide theoretical explanations for several previously unexplained phenomena in Park et al. (2024), including why ICL can suppress the original meaning of each word learned in pre-training (Section˜6.1), why the learned representations form an apparently regular yet slightly distorted structure (Sections˜6.2 and 6.3), and why the energy decreases but does not converge to zero (Section˜6.4);
@@
    Our theory highlights an implicit low-frequency bias in ICL, and predicts that LLMs are naturally robust to errors in the input prompts, which we have verified empirically (Section˜6.5).
This paper is organized as follows. In Section˜2, we review prior work on the mechanisms of in-context learning and the structure of hidden representations, emphasizing the core challenges and introducing core ideas of our approach. In Section˜3, we define notations used and the problems considered throughout the paper, as well as outline our method. In Section˜4, we present our results on context-wise convergence, and in Section˜5, we establish the main result by combining context-wise and layer-wise convergence. In Section˜6, we discuss the theoretical implications of our framework and address specific empirical observations. Finally, we conclude in Section˜7 with a summary and outlook for future work. Due to space constraints, we defer the formal proofs of our theoretical results and the empirical validation of our assumptions to the appendix.
@@ 3 Preliminaries
[…]
@@ 3.1 Data Generation
[…]
@@ 3.2 Model Architecture
Throughout the paper, we consider a simplified yet deep and non-linear Transformer model, as described in Algorithm˜1, where $d\in\mathbb{N}$ is the hidden and input dimension, $L\in\mathbb{N}$ is the number of layers, ${\bm{A}}^{(\ell)}=\left\{a_{k,j}^{(\ell)}\right\}_{k,j\in[n]}$ is the (single-head) attention map at layer $\ell$, and $\sigma^{(\ell)}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the neuron-wise transformations at layer $\ell$ (which may include feedforward networks (FFNs), normalization layers, and other non-linearities).
[…]
This model described in Algorithm˜1 also omits several other standard components such as normalizations and residual connections. We note that this is to not over-complicating the theoretical results while still capturing the core mechanisms and challenges of the model. There is flexibility in our analysis to include other components, but we choose to focus on a minimal version in the main paper to clearly highlight the key ideas behind our theoretical results. See Appendix˜E for further discussion on integrating additional components.
@@ 3.3 Methodology Outline
As noted before, we treat the attention maps in the Transformer as given, rather than dynamically generated from the hidden representations. Specifically, we assume the attention maps in the model are composed of a class of structured maps we refer to as nice attentions, formally defined in Definition˜3. These are attention maps whose connectivity patterns are determined by a function of the input tokens. This concept can be viewed as a generalization of the notion of induction heads (see eq.˜1), as illustrated by the structural similarity between eq.˜1 and eq.˜4.
@@ 4 Context-wise Convergence
Notice that in Definition˜3, the $k$-th row of the attention map $\{a_{k,j}\}_{j=1}^{n}$ is only well defined only there exists $j\in[k]$ such that $x_{j}\in f(x_{k})$ (otherwise all attention weights in this row are 0, violating the assumption that each row of ${\bm{A}}$ sums up to 111). This is guaranteed under our setup because we have explicitly set the first $c$ tokens to be a traversal over the entire vocabulary.
We are now ready to state the main theorem of this section. Notice that this theorem stated here depends on the DGP, as it relies on the distribution of $F_{x,k}$. However, it is possible to prove a weaker but more general version of this theorem that is independent of the DGP. See Appendix˜D for more details.
@@
There exists a constant $C>0$ that only depends on $\mathcal{G}$, and an event with probability at least 0.9990.9990.999, such that within this event, the following statement holds. Suppose ${\mathcal{V}}=\left\{{\bm{v}}_{k}\right\}_{k=1}^{n}\in\left(\mathbb{R}^{d}\right)^{n}$ is a good sequence converging to ${\mathcal{Z}}=\left\{{\bm{z}}_{x}\right\}_{x\in[c]}$ with parameter $\gamma$, and ${\bm{A}}=\left\{a_{k,j}\right\}_{k,j\in[n]}\in\mathbb{R}^{n\times n}$ is a nice attention map with parameter $\psi$ that reflects a function $f:[c]\to\mathbbm{2}^{[c]}$. Then ${\bm{A}}{\mathcal{V}}$ is a good sequence converging to
@@
Theorem˜1 shows that applying a nice attention map that reflects a function to a good sequence yields another good sequence, and that the attention operation effectively acts on the latent representations. In other words, attention maps of this kind preserve convergence and transform latent representations in a token-consistent way.
@@ 5 Layer-wise Convergence
In Theorem˜1, we showed how latent representations evolve under a single attention layer. In this section, we study how these latent representations change across layers.
@@
[…]
@@ 5.1 Main Results
There exists an event with probability at least 0.990.990.99 such that the following statement holds. Let $n_{[x]}$ be the largest $k\in[n]$ such that $x_{k}=x$. Let ${\bm{M}}={\bm{D}}^{-1/2}{\bm{W}}{\bm{D}}^{-1/2}$. Let the eigen-decomposition of ${\bm{M}}$ be
@@
where the eigenvalues are arranged in a non-increasing order of absolute values; ${\bm{\Lambda}}$ contains $q-1$ eigenvalues and ${\bm{\Lambda}}^{\prime}$ contains $n-q$ eigenvalues. Let $\bm{A}^{(\ell)}$ be defined as in eq.˜7, and $\left\{{\bm{v}}_{k}^{(\ell)}\right\}_{k=1}^{n}$ be defined as in Algorithm˜1. Suppose each $\sigma^{(\ell)}$ is a great mapping w.r.t. $\left(\gamma_{1}^{(\ell)},\gamma_{2}^{(\ell)},\mathbb{R}^{d},\bm{X}\right)$ and $\left(\gamma_{1}^{\prime(\ell)},\gamma_{2}^{\prime(\ell)},\mathbb{R}^{d},\bm{Y}\right)$, and have finite Lipschitz constant. Suppose there exists $\epsilon>0$ satisfying that $\delta_{q}\left(\rho_{A}^{(\ell)}{\bm{I}}+\rho_{B}^{(\ell)}{{\bm{M}}}\right)\frac{\gamma_{1}^{(\ell)}}{\gamma_{2}^{\prime(\ell)}}\leq 1-\epsilon$ for all $\ell\in[L]$. Then
@@ 6 Discussion About the Theoretical Results
In Theorem˜2, we proved that the representations converge to the top eigenspace of ${\bm{M}}={\bm{D}}^{-1/2}{\bm{W}}{\bm{D}}^{-1/2}$. This matrix has been extensively studied in the literature on graph learning and spectral methods (Kipf & Welling, 2016; Li et al., 2018; Wu et al., 2019; Yang et al., 2021; Spielman, 2019; Trevisan, 2013). Specifically, since ${\bm{W}}$ is a symmetric matrix with non-negative entries, it defines a (weighted) undirected graph. Let $\widehat{\bm{L}}$ be the symmetrically normalized Laplacian matrix of this graph (Spielman, 2019), it holds that $^{\bm{M}}={\bm{I}}-\widehat{\bm{L}}$, i.e. ${\bm{M}}$ and $\widehat{\bm{L}}$ share the same set of eigenvectors, with the order of eigenvalues being reversed. Thus, top eigenvectors of ${\bm{M}}$ corresponds to low eigenvalues of $\widehat{\bm{L}}$, which is known to encode the low-frequency (smooth) and low-energy signals on the graph, as they tends to assign similar values to adjacent nodes. These low eigenvectors of $\widehat{\bm{L}}$ are often used as coordinates for graph visualization, as it is know that they form figures that match human intuition (Tutte, 1963) [^2] Likely because humans also have a low-frequency bias in visual processing., and exactly explains why the ICLR phenomenon, where the hidden representations encode global graph structure, emerges in such settings.
We confirm this theoretical prediction by reproducing the experiments in Park et al. (2024) and compare the principal components of the actual hidden representations and the analytical prediction, i.e. top eigenvectors of ${\bm{W}}$. The result is shown in in Figure˜2. It is clear that the analytical predictions align closely with the empirical principle components [^3] All the experiments here and below are conducted with llama-3.1-8B using the NNsight library (Fiotto-Kaufman et al., 2025)..
@@ 6.2 Why Start From the 2nd Eigenvector?
In both Theorem˜2 and Figure˜2, we intentionally omit the first eigenvector of ${\bm{M}}$. This is due to the coincidental alignment between the Laplacian and PCA. In short, the 1st eigenvector of ${\bm{M}}$ corresponds to a constant vector added to each ${\bm{z}}_{x}^{\prime}$; however, PCA involves a centralization step that removes the mean component from the data. Specifically, it can proved that the first eigenvector of ${\bm{M}}$ is exactly ${\bm{d}}^{-1/2}$ (Spielman, 2019), and the centralization operation is to projecting ${\bm{V}}_{n}^{(L)}$ onto the space orthogonal to ${\bm{1}}$, which is equivalent to projecting ${\bm{V}}_{n}^{(L)}{\bm{D}}^{1/2}$ onto the space orthogonal to ${\bm{d}}^{-1/2}$, effectively removing the component aligned with the first eigenvector of ${\bm{M}}$.